# WixUp: A General Data Augmentation Framework for Wireless Perception in Tracking of Humans

### v1

@@ 5.1.3 Baseline augmentation methods
Baseline Null: no augmentation. Initially, we employ a baseline approach using solely the original real data without any augmentation.
@@
Baseline CGA: conventional global augmentation Since there is a lack of prior general DA method for all tasks and models discussed in this paper, we look into conventional global DA in general 2D and 3D data as our baselines instead. Therefore, we opt for a conventional global augmentation, named CGA, involving random scaling within the range of 0.8 to 1.2. Although it can not facilitate unsupervised domain adaptation through self-training due to its non-mixing-based approach, it is applicable in all the benchmarking experiments we will delve into in the following. More specifically, in pose estimation, the scaling is applied to both the sensing point clouds and the ground truth points. As for the other two tasks of user and action recognition, the scaling does not apply to the class labels.
@@
Baseline Stack: random duplication. The MiliPoint [17] dataset also introduces a simple DA technique, named stack. This method involves zero-padding each frame to standardize the number of points. Subsequently, they randomly resample from these points, serving as one data duplication. According to the original paper, this procedure is repeated several times for each frame: five times for tasks involving pose estimation and action recognition, and fifty times for the user identification task. Intuitively, this approach not only ensures a uniform input size but also enhances the data diversity through randomness. However, our reproduction results show that it actually negatively impacts the model performance in many cases. So, we only run it in initial benchmarking experiments and refrain from the rest.
@@
WixUp+: a combination of WixUp and CGA. Except for the self-training experiments, it is valid to use WixUp on top of Baseline CGA in benchmarking experiments. In practice, it is common to employ multiple data augmentation methods simultaneously in many machine learning applications. Therefore, we combine WixUp with CGA as another comparison with only WixUp. However, in our experiment results, we observed that the inclusion of CGA, noted as WixUp+, does not always bring additional performance enhancements, a point we will delve into in the forthcoming analysis section. Overall, WixUp itself constantly increases performance beyond the capabilities of individual baselines in most cases.
5.1.4 Datasets
@@
In our experiments, we leverage three mmWave datasets and one additional acoustic dataset for the cross-modality experiment. Specifically, they are three publicly available mmWave human tracking datasets alongside an acoustic hand-tracking dataset collected from our prior research. This diverse selection enables us to demonstrate the generalizability of WixUp across datasets and sensing modalities.
@@
MiliPoint Dataset, published recently in the NeurIPS dataset track [17], focuses on low-intensity cardio-burning fitness movements. With 49 distinct actions and a massive dataset of 545,000 frames from 11 subjects, it surpasses previous datasets in both action diversity and data volume. It also covers pose estimation and user identification. We use this dataset as the main test bed for the benchmarking experiments in this work. The data collection process considers factors such as movement intensity and diversity. Participants were asked to perform a series of movements while being monitored by a mmWave radar and a Zed 2 Stereo Camera for the ground truth. The Texas Instrument IWR1843 mmWave radar, a common choice for wireless perception research, operates between 77 GHz to 81 GHz with a chirp duration of 100 us and a slope of 40 MHz/us; it has a large bandwidth of 4GHz and thus can achieve a range resolution of 4 cm. The Zed 2 Stereo Camera complemented this by providing the ground-truth 3D skeleton of 18 keypoints, by getting a depth map from the disparity between two views and then feeding it into a neural network for pose estimation. This offers reliable accuracy for their settings in both 3 and 15 meters away. However, this public dataset consists solely of the 3D point clouds without the raw mmWave data as we discussed above, leaving around 8-22 points per frame. So, we show WixUp can inversely use the processed data for DA.
@@
MARS Dataset (Millimeter-wave Assistive Rehabilitation System) [7] is a pioneer work providing large-scale datasets in wireless sensing, designed for the rehabilitation of motor disorders utilizing mmWave. This work comes with a first-of-its-kind dataset of mmWave point cloud data, featuring 70 minutes of 10 different rehabilitation movements performed by 4 human subjects, providing 19 human keypoints and 40,083 labeled frames alongside video demonstrations made public. One common Texas Instrument IWR1443 Boost mmWave radar runs data acquisition at 76-81GHz. The chirp duration is 32us with a slope of 100MHz/us, facilitating a range resolution of 4.69cm and a maximum detection range of 3.37cm. Additionally, a Kinect V2 with infrared depth cameras captures ground truth, collocated with the radar and synchronized with the Kinect’s fixed sampling rate of 30 Hz. Similarly, their signal processing only keeps the first 64 points, consisting of a 5D time-series point. We take the subset of solely the coordinates to run the same experiment set up with the other mmWave dataset we use.
@@
MMFi Dataset stands out as another large-scale wireless dataset [53]. It features as the first multi-modal non-intrusive 4D human dataset, including mmWave radar, LiDAR, WiFi CSI, infrared cameras, and RGB cameras, emphasizing the fusion of sensors for multi-modal perception. The total has 320,000 synchronized frames across five modalities from 40 human subjects with 27 categories of daily and rehabilitation actions, providing a valuable view for both everyday and clinical research in human motion. The radar data is collected by a Texas Instrument IWR6843 60-64GHz mmWave. The detailed parameters for the chirp were not disclosed. Moreover, a novel mobile mini-PC captures and synchronizes data from multiple sensors, allowing for data collection in diverse environments. So, we use this dataset in our self-training experiments to test unsupervised domain adaptation across environments.
@@
Acoustic Dataset: We collected an acoustic dataset [1], which is a sensing platform designed for human hand tracking, employing the same FMCW (frequency-modulated continuous wave) signal algorithms utilized by mmWave radar systems. It features fine-grained continuous tracking of 21 highly self-occluded finger joints in 3D. Data collection involved 11 participants across three distinct environments, yielding a total of 64 minutes of meticulously selected hand motion data, covering a wide range of expressive finger joint movements. The hardware setup comprises a development microphone array board alongside a speaker and a Leap Motion infrared camera utilized solely for collecting ground truth during training. The 7-channel mic array shares the same layout and sensitivity specifications as the Amazon Echo 2 Home assistant. The system emits ultrasound modulated into 17k-20kHz FMCW chirps with a duration of 10ms. Subsequently, through the de-chirping algorithm, it achieves a range resolution of 3.57mm with only a small bandwidth, owing to the low speed of sound. This acoustic dataset helps demonstrate WixUp’s versatility across multiple wireless sensing modalities and its flexibility in handling other formats of input data.
5.1.5 Implementation
@@
To be able to run a wide range of experiments across datasets, tasks, model architectures, and baselines, we need a flexible code framework. Two of the three mmWave datasets, MARS and MMFi, do not come with a full code base, so we rewrite the code base of MiliPoint as a uniform framework to train and evaluate all the datasets. It supports a variety of model architectures including DGCNN, PointNet++, and PointTransformer, along with three tasks. We intend to release our framework as open source upon the acceptance of this paper, aiming to facilitate further research in DA within this domain.
@@
Table 1: Our code framework implementation can reproduce the results reported in the MiliPoint [17] dataset.
Table 1 illustrates our reproduction result of MiliPoint, wherein we run 50 epochs on the full data set for each of the three tasks using the DGCNN model. It runs less than 10 hours on a single RTX 8000 for each task. Our reproduction yields slightly better results for user identification and action recognition in terms of top-one classification accuracy in percentage, and it maintains a similar MLE for 9-point keypoint pose estimation. In conclusion, the reproduction result is verified to align with the reported results in the original paper. Besides, the errors here are usually higher than those in the following benchmarking because those are trained with a subset of data and five times fewer training epochs, in order to facilitate the extensive benchmarking experiments of various scenarios.
5.2 Benchmark the generalizability of WixUp in supervised learning
Following the above experiment setup, first, we assess how WixUp boosts model performance in supervised learning across diverse scenarios, encompassing datasets, model architectures, tasks, and sensing modalities. Broadly, WixUp expands the training data size for supervised learning, consistently delivering notable enhancements over the no-augmentation result and surpassing other baseline augmentation techniques.
5.2.1 Generalize across datasets
@@
Table 2: WixUp effectively reduces the error across diverse datasets and surpasses other baselines.
5.2.2 Generalize across model architectures
@@
Table 3: WixUp consistently improves performance across model DGCNN and PointTransformer, even outperforms CGA+WixUp sometimes owing to the instability of CGA.
The experimental results in Table. 3, highlight the performance across two distinct model architectures: DGCNN and Pointformer. And we run them for all three datasets of MiliPoint, MARS, and MMFi. The numbers in the table are errors of MLE in cm for keypoint pose estimation. As shown in the table, both our methods and the baselines bring model improvement across two model architectures. And WixUp consistently yields more improvements compared to Baseline CGA.
Moreover, the supplementary incorporation of WixUp alongside CGA, i.e., WixUp+, results in additional enhancements across most scenarios, except for PointTransformer on MARS and MMFi. Instead, WixUp performs even better without additional CGA. The incorporation of CGA, we think, doesn’t consistently result in improvements, possibly owing to the unstable nature of random scaling within CGA. The instability arises from the significant randomness inherent in random scaling, despite its straightforward usage in this context.
@@
In summary, WixUp is robust and might even outperform its combination of CGA (WixUp+) because it is a stable method. Overall, this table of results demonstrates its flexibility as a general module that can be integrated into downstream research works regardless of their modeling methods.
5.2.3 Generalize across human tracking tasks
@@
In brief, WixUp outperforms Baseline CGA in most cases. Furthermore, WixUp+, the combination of WixUp with CGA, leads to further improvements. Notably, while action recognition with WixUp falls short of CGA, WixUp+ surpasses CGA by a significant margin, reaching 84% of improvement. The reason could be that the overall accuracy in action recognition is low, so the percentage might diverge dramatically.
@@
Table 4: WixUp generalize well across three tasks.
@@
Figure 4: Self-training results for unsupervised domain adaptation across users and environments.
5.2.4 Generalize across sensing modality&data format
@@
In order to assess WixUp with different wireless sensing modalities and varied formats of raw data, we test with this acoustic dataset along with its dedicated code base using a CNN+LSTM model. As outlined above 5.1.4, this dataset entails an acoustic sensing system designed for tracking 21 finger joints in 3D. To apply WixUp with acoustic range profiles, we simply bypass the step of simulating from coordinates to range profiles. Instead, we directly intersect the range profile to generate new range profiles as model input. In contrast, the Baseline CGA does not directly apply to the range profile anymore. Instead, we keep the straightforward data augmentation method by slightly shifting the range profile along the range axis; and we use the original accuracy reported in the paper as a baseline comparison with WixUp.
In summary, the original pose estimation yielded a mean absolute error (MAE) of 13.93mm for user-dependent testing. After running the same user-dependent test with WixUp, we achieved an MAE of 10.56mm, which closely approaches the best results reported by the paper achieved by user-adaptive testing.
@@
5.3 WixUp reduces labeling efforts by unsupervised domain adaptation via self-training
@@
Beyond testing our data augmentation in standard supervised learning for wireless perception, we extend our objective to show its capability of self-training, founded on its scheme of mixing data to synthesize new data. In this subsection, we conduct another group of experiments that illustrates WixUp’s capability to significantly reduce the effort for labeling data by utilizing unsupervised domain adaptation (UDA) through self-learning.
@@
Moreover, it is worth mentioning that non-mix-up methods like CGA and stacking can NOT be applied to UDA for self-training. Because they augment one data point at a time instead of mixing two data points from two distributions. In contrast, synthesizing data by mixing two points allows for the creation of more diverse and realistic samples, aiding in domain adaptation via self-training.
@@
In practice, this self-training process proves especially advantageous in wireless sensing, where acquiring new data often demands significant manual efforts, and system performance might significantly drop in new rooms or new users.
5.3.1 Unsupervised domain adaptation across users

### v2

@@ 5.1.3. Baseline augmentation methods
Baseline Null: no augmentation. One baseline approach uses solely the original real data without any augmentation.
@@
Baseline CGA: conventional global augmentation Due to the lack of prior generic DA for all tasks and models discussed in this paper, we look into conventional global DA in general 2D and 3D data as our baselines. Therefore, we opt for a conventional global augmentation, named CGA, involving random scaling within the range of 0.8 to 1.2. Although it can not facilitate unsupervised domain adaptation through self-training due to its non-mixing-based approach, it is applicable in all the benchmarking experiments we will delve into in the following. Specifically, in pose estimation, scaling is applied to both sensing point clouds and ground truth points, while for user and action recognition, scaling is not applied to class labels.
@@
Baseline Stack: random duplication. The MiliPoint (cui2024milipoint, ) dataset also introduces a simple DA technique, named stack. This method involves zero-padding each frame to standardize the number of points. Subsequently, they randomly resample from these points, serving as one data duplication. According to the original paper, this procedure is repeated several times for each frame: five times for tasks involving pose estimation and action recognition, and fifty times for the user identification task. Intuitively, this approach not only ensures a uniform input size but also enhances the data diversity through randomness. However, our reproduction results show that it actually negatively impacts the model performance in many cases. So, we only run it in initial benchmarking experiments and refrain from the rest.
@@
WixUp+: a combination of WixUp and CGA. Except for the self-training experiments, it is valid to use WixUp on top of Baseline CGA in benchmarking experiments. In practice, it is common to employ multiple DA methods simultaneously in many machine-learning applications. Therefore, we combine WixUp with CGA as another comparison with only WixUp. However, in our experiment results, we observed that the inclusion of CGA, noted as WixUp+, does not always bring additional performance enhancements, a point we will delve into in the forthcoming analysis section. Overall, WixUp itself constantly increases performance beyond the capabilities of individual baselines in most cases.
5.1.4. Datasets
@@
In our experiments, we leverage three mmWave datasets and one additional acoustic dataset for the cross-modality experiment. Specifically, they are three publicly available mmWave human tracking datasets alongside an acoustic hand-tracking dataset collected from our prior research. This diverse selection could demonstrate the generalizability of WixUp across datasets and sensing modalities.
@@
MiliPoint Dataset (cui2024milipoint, ) focuses on low-intensity cardio-burning fitness movements. With 49 distinct actions and a massive dataset of 545,000 frames from 11 subjects, it surpasses previous datasets in both action diversity and data volume. It also covers pose estimation and user identification. We use this dataset as the primary test bed for our benchmarking experiments. The data collection process considers factors such as movement intensity and diversity. Participants perform a series of movements while monitored by a mmWave radar and a Zed 2 Stereo Camera for the ground truth, at a frame rate of around 24Hz. The Texas Instrument IWR1843 mmWave radar, a common choice for wireless perception research, operates between 77 GHz to 81 GHz with a chirp duration of 100 us and a slope of 40 MHz/us, with a large bandwidth of 4GHz for a 4cm range resolution. The Stereo Camera provides ground truth of 18 3D keypoints, offering reliable accuracy for their settings at both 3 and 15 meters away. However, this dataset consists solely of 3D point clouds without the raw mmWave data, leaving around 8-22 points per frame. So, we could show WixUp can inversely use the processed data for DA.
@@
MARS Dataset (Millimeter-wave Assistive Rehabilitation System) (marsdataset2021, ) is a pioneer work providing large-scale datasets in wireless sensing, designed for the rehabilitation of motor disorders utilizing mmWave. This work comes with a first-of-its-kind dataset of mmWave point cloud data, featuring 70 minutes of 10 different rehabilitation movements performed by 4 human subjects, providing 19 human keypoints and 40,083 labeled frames alongside video demonstrations made public. One common Texas Instrument IWR1443 Boost mmWave radar runs data acquisition at 76-81GHz. The chirp duration is 32us with a slope of 100MHz/us, facilitating a range resolution of 4.69cm and a maximum detection range of 3.37cm. Additionally, a Kinect V2 with infrared depth cameras captures ground truth, collocated with the radar and synchronized with the Kinect’s fixed sampling rate of 30 Hz. Similarly, their signal processing only keeps the first 64 points, consisting of a 5D time-series point. We take the subset of solely the coordinates to run the same experiment set up with the other mmWave dataset we use.
@@
MMFi Dataset stands out as another large-scale wireless dataset (yang2024mm, ), featuring the first multi-modal non-intrusive 4D human dataset, including mmWave radar, LiDAR, WiFi CSI, infrared cameras, and RGB cameras, for the fusion of sensors and multi-modal perception. The total has 320,000 synchronized frames across five modalities from 40 human subjects with 27 categories of daily and rehabilitation actions, providing a valuable view for both everyday and clinical research in human motion. The radar is Texas Instrument IWR6843 60-64GHz mmWave, whose detailed parameters were not disclosed. The RGB and depth cameras both have a frame rate of 30Hz and the radar is set correspondingly. Moreover, a novel mobile mini-PC allows data collection in diverse environments. So, we use this dataset in our self-training experiments to test unsupervised domain adaptation across environments.
@@
Acoustic Dataset: We collected an acoustic dataset in  (acousticdataset2023, ), which is a sensing platform for human hand tracking, employing the same FMCW signal algorithms of mmWave radars. It features fine-grained continuous tracking of 21 highly self-occluded finger joints in 3D. Data was collected from 11 participants across three environments, totaling 64 minutes of selected hand motions, covering expressive finger joint movements. The hardware comprises a development microphone array board, a speaker, and a Leap Motion infrared camera for ground truth (frame rate is 90$\sim$110Hz). The 7-channel mic array shares the same layout and sensitivity specifications as Amazon Echo 2. The system emits ultrasound modulated into 17k-20kHz chirps with a 10ms duration, achieving a range resolution of 3.57mm, owing to the low speed of sound and the custom de-chirping. This dataset helps demonstrate WixUp’s versatility across multiple wireless modalities and flexibility in handling other input data formats.
Note that the wireless dataset’s frame rate is usually limited to its ground truth camera’s frame rate, which is typically around 30Hz. As the chirp duration reported above, the chirp rate with no idle time could be up to a few KHz for mmWave sensing, and 100Hz for acoustic sensing. While the wireless signal data could have a higher frame rate, the experiment settings are aligned at a similar scale with ground truth camera’s frame rate. After all, one advantage of our method is that it is designed to work in the absence of this kind of parameter.
5.1.5. Implementation
@@
To run a wide range of experiments across datasets, tasks, model architectures, and baselines, we need a flexible code framework. Since, two of the three mmWave datasets, MARS and MMFi, do not come with a full code base, we rewrite the codebase of MiliPoint as a uniform framework for testing all datasets. It supports a variety of model architectures including DGCNN, PointNet++, and PointTransformer, along with three tasks.We published our code of WixUp framework on Github [^1] https://github.com/lydhr/WixUp, aiming to facilitate further research in DA within this domain.
@@
Table 1. Our code base implementation can reproduce the results reported in the MiliPoint (cui2024milipoint, ) dataset.
Table 1 illustrates our reproduction result of MiliPoint, wherein we run 50 epochs on the full data set for each of the three tasks using the DGCNN model. It runs less than 10 hours on a single RTX 8000 for each task. Our reproduction yields slightly better results for user identification and action recognition in terms of top-one classification accuracy, and it maintains a similar MLE for 9-point keypoint pose estimation. In conclusion, the reproduction result is verified to align with the reported results in the original paper. Besides, the errors here are usually higher than those in the following benchmarking because those are trained with a subset of data and five times fewer training epochs, to facilitate the extensive benchmarking experiments of various scenarios.
5.2. Benchmark the generalizability of WixUp in supervised learning
Following the above setup, first, we assess how WixUp boosts model performance in supervised learning across diverse scenarios, encompassing datasets, model architectures, tasks, and sensing modalities. Broadly, WixUp expands the training data size for supervised learning, consistently delivering notable enhancements over the no-augmentation result, surpassing other baseline augmentations.
5.2.1. Generalize across datasets
@@
Table 2. WixUp effectively reduces the error across diverse datasets and surpasses other baselines.
5.2.2. Generalize across model architectures
@@
Table 3. WixUp consistently improves performance across model DGCNN and PointTransformer, even outperforms CGA+WixUp sometimes owing to the instability of CGA.
The experimental results in Table. 3 highlight the performance across two distinct model architectures: DGCNN and Pointformer, tested on all three datasets of MiliPoint, MARS, and MMFi. The numbers in the table are MLE errors in cm for keypoint pose estimation. In conclusion, WixUp consistently yields more improvements compared to Baseline CGA. Moreover, the supplementary incorporation of WixUp alongside CGA (WixUp+) brings additional enhancements across most scenarios, except for PointTransformer on MARS and MMFi where WixUp along performs better. The incorporation of CGA, we think, doesn’t consistently result in improvements, possibly owing to the unstable nature of random scaling within CGA. The instability arises from the significant randomness inherent in random scaling, despite its straightforward usage in this context.
@@
In summary, WixUp is robust and may outperform its combination of CGA (WixUp+) due to its stability. Overall, this table of results highlights its flexibility as a general module, applicable to downstream tasks regardless of their methods.
5.2.3. Generalize across human tracking tasks
@@
In brief, WixUp outperforms Baseline CGA in most cases. Furthermore, WixUp+, the combination of WixUp with CGA, leads to further improvements. Notably, while action recognition with WixUp falls short of CGA, WixUp+ surpasses CGA by a significant margin, reaching 84% of improvement. The reason could be that the overall accuracy in action recognition is low, so the percentage might diverge dramatically. Moreover, expanding the augment data size by tweaking WixUp brings further drops in errors, which we will ablate study in the following section. Besides, WixUp is a mixing-based local augmentation, providing more flexibility than global augmentation like CGA.
@@
Table 4. WixUp generalize well across three tasks.
@@
Figure 4. Self-training results for unsupervised domain adaptation across users and environments.
5.2.4. Generalize across sensing modality&data format
@@
To assess WixUp with different wireless sensing modalities and raw data formats, we test with this acoustic dataset along with its dedicated code base using a CNN+LSTM model. As outlined in  5.1.4, this dataset entails an acoustic sensing system designed for tracking 21 finger joints in 3D. To apply WixUp with acoustic range profiles, we simply bypass the step of simulating from coordinates to range profiles. Instead, we directly intersect the range profile to generate new range profiles as model input. In contrast, the Baseline CGA does not directly apply to the range profile anymore. So, we keep the straightforward DA method by slightly shifting the range profile along the range axis; and we use the original accuracy reported in the paper as a baseline comparison with WixUp.
In summary, the original mean absolute error (MAE) is 13.93mm for user-dependent pose estimation. By running the same test via WixUp, we achieved a better MAE of 10.56mm(+24.19%), which closely approaches the best results reported by the paper achieved by user-adaptive testing. In this regard, we observe that generalizing across modalities should be less challenging than that across data formats. Theoretically, this could be attributed to the fact that the primary distinction between modalities lies in signal frequency. When processed by the standard pipeline in Fig. 2, different modalities’ data exhibits similar characteristics at the range profile level, which are generally a mixture of Gaussians as visualized in Fig. 3.
@@
5.3. WixUp reduces labeling efforts by unsupervised domain adaptation via self-training
@@
Beyond testing WixUp in standard supervised learning, we extend our objective to show its capability of self-training, founded on its scheme of mixing data to synthesize new data. In this subsection, another group of experiments illustrates WixUp’s capability to significantly reduce the effort for labeling data by utilizing unsupervised domain adaptation (UDA) through self-training.
@@
Moreover, it is worth mentioning that non-mix-up methods like CGA and stacking can NOT be applied to UDA for self-training. Because they augment one data point at a time instead of mixing two data points from two distributions. In contrast, synthesizing semi-real labeled data by mixing two points allows for the creation of more diverse and realistic samples, aiding in domain adaptation via self-training.
@@
In practice, this self-training process is particularly advantageous in wireless sensing, where acquiring new data often demands significant manual efforts, and system performance might drop in new users or environments.
To clarify, domain adaptation is a broad topic encompassing a wide range of approaches beyond data augmentation, such as adversarial training, contrastive learning, feature transformation, etc. which are beyond the scope of this paper. This work primarily focuses on proposing a data augmentation framework for wireless sensing, with domain adaptation being one of its usages.
5.3.1. Unsupervised domain adaptation across users